OpenStudy (anonymous):

Please help me, medal/fan I DONT WANT AN ANSWER A volleyball playoff competition has eliminations each round. The table below shows the number of teams in each round of the tournament: Round (x) 1 2 3 4 Teams f(x) 16 8 4 2 Compute the average rate of change of f(x) from x = 1 to x = 4 and identify the meaning of that rate.

OpenStudy (anonymous):

Answers: . negative three over fourteen; on average, there were 3 fewer teams every 14 rounds negative fourteen over three; on average, there were 14 fewer teams every 3 rounds 3; there were 3 rounds between rounds 1 and 4 14; there were 14 fewer team between rounds 1 and 4,

Vocaloid (vocaloid):

average rate of change from x = 1 to x = 4 average rate of change = [f(4)-f(1)]/(4-1)

Vocaloid (vocaloid):

first step: find f(4) (what is the value of the function when x = 4?)

OpenStudy (anonymous):

f(4)-f(1)/3 @Vocaloid

Mehek (mehek14):

(1,16) and (4,2) \(\dfrac{y_2-y_1}{x_2-x_1}\\ y_2=2\\y_1=16\\x_2=4\\x_1=1\)

Mehek (mehek14):

so \(\dfrac{2-16}{4-1}\)

Mehek (mehek14):

what is 2 - 16 and 4 - 1 ?

Mehek (mehek14):

@Clueless_much78 ?

OpenStudy (anonymous):

-14 and 3 @Mehek14 sorry family dinner :(

Mehek (mehek14):

ok so \(\dfrac{-14}{3}\) is that in your answer choices?

OpenStudy (anonymous):

yes, it is!! thanks!! :)